### v2

# Agentic Robot: A Brain-Inspired Framework for Vision-Language-Action Models in Embodied Agents

Abstract
Long-horizon robotic manipulation poses significant challenges for autonomous systems, requiring extended reasoning, precise execution, and robust error recovery across complex sequential tasks. Current approaches, whether based on static planning or end-to-end visuomotor policies, suffer from error accumulation and lack effective verification mechanisms during execution, limiting their reliability in real-world scenarios. We present Agentic Robot, a brain-inspired framework that addresses these limitations through Standardized Action Procedure (SAP)–a novel coordination protocol governing component interactions throughout manipulation tasks. Drawing inspiration from Standardized Operating Procedures (SOPs) in human organizations, SAP establishes structured workflows for planning, execution, and verification phases. Our architecture comprises three specialized components: (1) a large reasoning model that decomposes high-level instructions into semantically coherent subgoals, (2) a vision-language-action executor that generates continuous control commands from real-time visual inputs, and (3) a temporal verifier that enables autonomous progression and error recovery through introspective assessment. This SAP-driven closed-loop design supports dynamic self-verification without external supervision. On the LIBERO benchmark, Agentic Robot achieves state-of-the-art performance with an average success rate of 79.6%, outperforming SpatialVLA by 6.1% and OpenVLA by 7.4% on long-horizon tasks. These results demonstrate that SAP-driven coordination between specialized components enhances both performance and interpretability in sequential manipulation, suggesting significant potential for reliable autonomous systems. Project Github: https://agentic-robot.github.io.
1 Introduction
Recent advances in foundation models have demonstrated remarkable potential for creating embodied agents capable of interpreting natural language instructions and executing complex manipulation tasks [3, 16, 14, 6, 2]. These systems effectively bridge the gap between high-level reasoning and low-level physical control. However, existing embodied manipulation systems struggle to achieve reliable performance on long-horizon tasks that require extended sequences of coordinated actions [11, 13, 7]. Real-world scenarios such as table setting, grocery packing, or furniture assembly demand not only sophisticated reasoning and precise motor control, but also robust error detection and recovery mechanisms throughout extended task execution [40, 4].
Through extensive analysis of current approaches, we identify fundamental limitations that prevent reliable long-horizon manipulation. Most existing methods fall into two categories with critical weaknesses: static plan-following agents that generate fixed execution sequences without adaptive feedback [16, 3], and end-to-end visuomotor policies that directly map observations to actions without intermediate reasoning [14]. Static planners suffer from compounding error propagation–small deviations early in execution cascade into catastrophic failures [32]. End-to-end policies lack mechanisms for introspection and often fail to recover from unexpected states, particularly when encountering scenarios outside their training distribution [40].
Drawing insights from Standardized Operating Procedures (SOPs) in human organizations [9, 31], we observe that reliable task execution requires structured coordination protocols. In natural cognition, complex behaviors emerge from specialized neural circuits working through well-defined interaction patterns: prefrontal regions handle planning, motor cortices execute actions, and sensory-motor loops provide continuous verification feedback [26, 21]. Similarly, in human organizations, SOPs establish clear workflows that minimize errors and enable effective collaboration across different roles. This biological and organizational wisdom suggests that robotic systems can benefit from structured coordination protocols that govern component interactions.
Inspired by these insights, we design Agentic Robot, a brain-inspired framework that introduces Standardized Action Procedure (SAP)–a novel coordination protocol specifically designed for embodied manipulation tasks. Unlike SOPs, which govern human workflows, SAP encodes the natural cognitive cycle into structured agent interactions for robotic systems. SAP defines the complete agentic loop that governs how our three specialized components–Planner, Executor, and Verifier–coordinate throughout task execution through well-defined interfaces and standardized protocols for information exchange, progress monitoring, and error recovery. Besides, Agentic Robot requires agents to maintain structured interaction protocols throughout the manipulation process. SAP ensures that task decomposition, action execution, and progress verification follow consistent procedure, dramatically reducing error accumulation while enabling robust recovery from failures. More specifically, all components follow strict SAP-defined workflows, ensuring that information handoffs comply with established protocols and eliminating the communication breakdowns that plague existing systems.
Our main contributions are as follows:
We introduce Agentic Robot, a brain-inspired agentic framework for embodied manipulation that incorporates structured coordination protocols. The framework is highly modular and interpretable, with well-defined component interfaces, making it a powerful platform for developing reliable long-horizon manipulation systems.
We propose Standardized Action Procedure (SAP), a novel coordination protocol that governs the complete agentic loop in robotic manipulation tasks. SAP encodes structured interactions between planning, execution, and verification phases, enhancing system reliability and reducing error propagation through standardized workflow management.
We achieve state-of-the-art performance on the LIBERO benchmark with an average success rate of 79.6%. Extensive experimental results convincingly demonstrate that our SAP-driven approach represents a promising framework for reliable embodied manipulation, with particularly strong improvements on challenging long-horizon tasks.
2 Agentic Robot Framework: A Brain-Inspired Control Loop
2.1 Overview
We introduce Agentic Robot, an agentic framework that reformulates long-horizon manipulation as a closed perception-reasoning-execution-verification loop, inspired by biological cognition and multi-agent LLM systems [9, 31]. Drawing insights from SOPs that govern effective human workflows, we propose SAP–a novel coordination protocol that structures component interactions throughout the manipulation process. SAP establishes explicit protocols for information exchange, progress monitoring, and error recovery, enabling robust execution of complex manipulation tasks. Our design is grounded in recent advances across large reasoning models (LRMs), vision-language models (VLMs), and vision-language-action (VLA) systems. We provide a detailed review of these foundations in the Related Works section (see Appendix B).
Our architecture integrates three specialized components: (1) a planner based on LRM that decomposes high-level instructions into structured subgoals, (2) an executor based on VLA that generates continuous control actions from subgoals and visual input, and (3) a verifier based on VLM that conducts self-assessment for autonomous progression or recovery. Each component operates within the SAP framework, following standardized interfaces and communication protocols that ensure seamless coordination throughout task execution.
As shown in Fig. 1, our agent processes task descriptions and RGB observations from third-person and egocentric cameras. The planner generates subgoals following SAP specifications, which the VLA model translates into 7-DoF actions based on visual input. Simultaneously, the verifier monitors a temporal frame buffer to determine subgoal completion according to SAP verification protocols, moving to the next subgoal upon success or triggering standardized recovery actions upon failure. This architecture implements a sequence of agentic steps, each combining intention grounding, visuomotor execution, and perception-based validation within the SAP framework, enabling execution correction without external supervision.
2.2 Planner: LRM for Subgoal Generation
The planner module, denoted as , functions as the high-level reasoning component within our SAP framework. It converts task instructions into a structured sequence of executable subgoals following standardized decomposition protocols:
where represents the initial visual observation. Each subgoal forms a complete and constrained instruction derived from an Atomic Skill Library [15], which defines standardized action templates such as:
pick up [object] | place [object] in/on [location] | turn on/off [device]
This constrained approach ensures compatibility with the executor while maintaining interpretability of the execution pipeline, adhering to SAP principles of structured component interaction.
We implement the planner using a state-of-the-art large multimodal reasoning model (e.g., GPT-4o), which processes both the instruction and optionally an image for visual grounding. The SAP-compliant prompt architecture includes three structured components: (1) a task preamble explaining the planner’s role within the framework, (2) the complete Atomic Skill Library specifying allowed action types, and (3) carefully selected few-shot examples demonstrating proper subgoal decomposition. These examples guide the model to establish appropriate task boundaries, resolve ambiguities, and break down complex instructions into 2-5 atomic steps. Through extensive validation, we determined that subgoals with 1-2 semantic units (e.g., verb + object or verb + object + location) achieve optimal balance between clarity and executability within the SAP framework.
2.3 VLA Executor: Reactive Visuomotor Policy
The executor module serves as the core visuomotor interface that transforms each subgoal and associated visual observations into continuous low-level control signals according to SAP execution protocols:
where represents the robot’s Cartesian displacement and gripper configuration. The first six dimensions encode translation and rotation vectors, while the last component indicates the binary gripper state.
We utilize OpenVLA [14], an open-source pretrained VLA model that establishes direct connections between natural language subgoals and visual observations. The architecture incorporates a large language model backbone with a visual transformer (ViT) to process multimodal inputs and generate appropriate motor commands. Each subgoal adheres to the structured format outlined in our Atomic Skill Library, enabling the VLA model to systematically generate actions by understanding both language instructions and visual scene content. This structured approach enhances compatibility and interpretability across manipulation scenarios while constraining the action space to physically feasible trajectories.
Despite its stateless design, the executor integrates robust error-handling capabilities through the SAP verification loop. When execution failures occur, the standardized verification mechanism detects issues through visual assessment and triggers specific recovery actions following SAP protocols. If multiple recovery attempts fail, the framework marks the task as failed and halts execution to prevent unsafe behaviors. This closed-loop error detection significantly improves system robustness–particularly in long-horizon tasks–by reducing cascading errors. As shown in Section 3.3, our method achieves up to 24% improvement over OpenVLA on the Bowl-Drawer task and 21% on Soup-Sauce, confirming the effectiveness of subgoal-level verification and recovery.
2.4 Verifier: Perception-based Subgoal Assessment and Recovery
The verifier module provides critical feedback within the SAP framework by assessing the success of each subgoal through visual analysis. For every verification step , it produces a binary response following a two-stage assessment protocol:
where is a sliding buffer of recent image pairs from third-person and wrist-mounted views. This temporal buffer captures visual dynamics such as object displacement or contact transitions, typically with and frame intervals of 5.
We employ Qwen2.5-VL-3B-Instruct [1] as the verifier model to evaluate whether subgoal is complete. The verification prompt follows SAP’s structured format: “Based on the image sequence, has the robot successfully completed [subgoal]?” The model is fine-tuned with LoRA [10] on a dataset of annotated triplets with . To adapt the verifier to subgoal-level introspection, we fine-tune Qwen2.5-VL-3B-Instruct using LoRA on a compact dataset of approximately 500 annotated triplets. Despite its small scale, the dataset covers diverse subgoal types and visual scenes, and leverages strongly structured prompts to guide learning. This setting demonstrates that even with limited supervision, targeted adaptation can yield effective visual verification in closed-loop execution (see Section 3.5).
When the initial response is , the verifier performs a secondary check to determine whether the robot is stuck:
where is a diagnosis module that detects conditions such as stationary arms, failed grasp, or oscillating behaviors. If , a recovery action is triggered:
such as lifting the gripper or reorienting the wrist. The system then re-executes and resumes the same two-stage verification process at the next interval. After unsuccessful recovery attempts, the task is marked as failed.
To optimize responsiveness and efficiency, verification is performed every 20 frames (i.e., Hz), achieving near-optimal accuracy (only 1.2% drop from 10-frame intervals) while reducing computational load by 48%. Compared to single-pass goal-checking methods, our two-level verifier allows for mid-execution correction and fine-grained failure localization.
2.5 SAP: Standardized Action Procedure for Coordinated Agentic Control
In robotic manipulation, the absence of structured coordination protocols often leads to execution failures, particularly in long-horizon tasks where accumulated errors and lack of systematic verification result in task breakdown. Drawing inspiration from Standardized Operating Procedures (SOPs) that have proven effective in complex collaborative environments, we introduce Standardized Action Procedure (SAP) as a systematic framework that encodes proven coordination patterns into robotic agentic systems.
SAP represents a principled approach to orchestrating closed-loop execution within our agentic robot framework by establishing standardized coordination protocols across perception, planning, execution, and verification components. The core design philosophy rests on three fundamental principles: (1) Modular Decomposition - complex manipulation tasks are systematically decomposed into manageable, verifiable subgoals as illustrated in Fig. 2; (2) Structured Coordination - component interactions follow predefined workflows rather than opportunistic communication; and (3) Adaptive Verification - systematic checkpoints enable early error detection and recovery.
2.5.1 SAP Operational Framework
Each SAP cycle at time constitutes an agentic step that encapsulates the complete perception-planning-execution-verification workflow:
where denotes egocentric and third-person views, represents the current subgoal within the structured task decomposition, is the executed action, and indicates the verification result.
SAP defines four specialized components with standardized interfaces and coordination protocols:
(1) Multimodal Perception. At each time step, the agent collects dual-perspective observations:
which provides comprehensive workspace understanding following standardized observation protocols.
(2) Formulated Plan. The planner converts task instructions into a structured sequence of executable subgoals following standardized decomposition protocols:
where represents the initial visual observation and each subgoal is derived from the Atomic Skill Library.
(3) Reactive Execution. The executor translates subgoal into low-level control signals:
where maps semantic goals and current vision into 7-DoF actions following standardized execution protocols.
(4) Temporal Verification. Every frames (typically 20), the verifier performs systematic evaluation:
where represents the sliding buffer of recent image pairs. If , the agent proceeds to the next subgoal. If not, and , a recovery action is triggered:
SAP execution is managed by an asynchronous finite-state machine with component-specific frequencies: the executor operates at 10 Hz (s), and the verifier at 0.5 Hz (s). By enforcing structured control cycles with modular boundaries and layered feedback, SAP enhances agent reliability and interpretability. It supports in-situ correction, isolates errors, and ensures safe recovery, addressing core limitations of open-loop or end-to-end systems in dynamic and uncertain manipulation environments.
3 Experiments
3.1 Experimental Setup
We evaluate our Agentic Robot framework on long-horizon manipulation tasks in simulated embodied environments. The agent employs a dual-camera perception system: a static agent-view camera for global scene context and a wrist-mounted eye-in-hand camera for local detail. Both cameras provide synchronized RGB observations at each timestep. The action space consists of a 7-dimensional continuous control vector representing 6-DoF end-effector control plus a binary gripper state.
Benchmarks. We carry out evaluations using the LIBERO benchmark suite [17], which provides a standardized way to assess instruction-following manipulation across various environments. Our experiments concentrate on four specific challenge subsets: LIBERO-Spatial, which focuses on understanding spatial relationships; LIBERO-Object, which tests generalization to new objects; LIBERO-Goal, which assesses abstract goal execution; and LIBERO-Long, which involves extended sequential manipulations. Each subset consists of 10 distinct tasks, and for each task, there are 50 human-teleoperated demonstrations.
Baselines. We benchmark our approach against the following generalist policies, including state-of-the-art open-sourced models: Diffusion Policy [5], Octo-Base [28], OpenVLA [14], TraceVLA [38], and SpatialVLA [23]. These methods represent a variety of model paradigms, including diffusion-based control (Diffusion Policy), transformer-based visuomotor policies (Octo-Base), and large-scale vision-language-action models (OpenVLA, TraceVLA, SpatialVLA). For fair comparison, we follow the original hyperparameters and evaluation settings as reported in their respective works without additional tuning. Detailed information are provided in Appendix A.2.
Implementation. Agentic Robot integrates the three modules described in Section 2: a GPT-4o-based planner for subgoal decomposition, an OpenVLA-based executor for visuomotor control, and a fine-tuned Qwen2.5-VL-3B-Instruct verifier for subgoal completion assessment. For error recovery, we raise the gripper to a safe position upon failure detection before re-evaluation. Unless otherwise specified, verification is performed every 20 frames.
3.2 Main Results
| Method | LIBERO-Spatial | LIBERO-Object | LIBERO-Goal | LIBERO-Long | Average |
| SR (Rank) | SR (Rank) | SR (Rank) | SR (Rank) | SR (Rank) | |
| Diffusion Policy [5] | 78.3 ± 1.1 (6) | 92.5 ± 0.7 (1) | 68.3 ± 1.2 (6) | 50.5 ± 1.3 (6) | 72.4 ± 0.7 (6) |
| Octo-Base (FT) [28] | 78.9 ± 1.0 (5) | 85.7 ± 0.9 (5) | 84.6 ± 0.9 (1) | 51.1 ± 1.3 (5) | 75.1 ± 0.6 (4) |
| OpenVLA (FT) [14] | 84.7 ± 0.9 (3) | 88.4 ± 0.8 (4) | 79.2 ± 1.0 (3) | 53.7 ± 1.3 (4) | 76.5 ± 0.7 (3) |
| TraceVLA (FT) [38] | 84.6 ± 0.2 (4) | 85.2 ± 0.4 (6) | 75.1 ± 0.3 (5) | 54.1 ± 1.0 (3) | 74.8 ± 0.5 (5) |
| SpatialVLA (FT) [23] | 88.2 ± 0.5 (1) | 89.9 ± 0.7 (2) | 78.6 ± 0.6 (4) | 55.5 ± 1.0 (2) | 78.1 ± 0.7 (2) |
| Agentic Robot (Ours) | 85.8 ± 0.6 (2) | 89.0 ± 0.8 (3) | 81.8 ± 0.8 (2) | 61.6 ± 1.2 (1) | 79.6 ± 0.8 (1) |
Table 1 presents success rates across four LIBERO benchmark suites. Agentic Robot achieves state-of-the-art performance with 79.6% average success rate, surpassing all baselines across diverse manipulation scenarios.
Cross-domain generalization. Agentic Robot consistently ranks among the top-3 performers across all task categories, demonstrating exceptional versatility in diverse manipulation scenarios. Unlike specialized approaches that exhibit domain-specific excellence but inconsistent cross-domain performance, our framework maintains balanced efficacy throughout the benchmark suite. This strong generalization indicates that our architecture effectively captures essential manipulation principles that go beyond task-specific requirements, which is a critical capability for deployment in unconstrained real-world environments.
Long-horizon planning. In the particularly challenging LIBERO-Long tasks, Agentic Robot significantly outperforms all baseline methods, achieving a 6.1% improvement over SpatialVLA, the previous state-of-the-art. This substantial enhancement is a direct result of our core architectural innovation: the breakdown of complex instructions into individually verifiable subgoals with clearly defined intermediate checkpoints. By implementing closed-loop verification at the subgoal level, our system effectively reduces error accumulation, which is a major limitation of existing approaches, especially as task complexity increases. The relationship between task horizon length and performance advantage is further examined in Section 3.3, where we demonstrate that Agentic Robot’s performance advantage grows in proportion to task complexity.
Verification and recovery. A key innovation in Agentic Robot is its explicit subgoal-level verification and recovery mechanism. Unlike end-to-end baselines relying on implicit success estimation, our system provides transparent execution monitoring with interpretable assessments, enabling targeted recovery strategies. This retry mechanism enhances performance across all benchmarks and proves particularly effective in complex multi-step tasks, which is illustrated in Fig. 3. Without recovery, baselines such as OpenVLA continue executing the current subgoal indefinitely, even when the gripper is clearly stuck, for instance, when it presses against the table after a failed grasp. In contrast, Agentic Robot uses visual verification to detect such failure states and triggers a simple recovery behavior: lifting the gripper vertically before retrying the action. As shown in the figure, this enables the robot to reattempt grasping and proceed to the next subgoal successfully. Although our current recovery policy is deliberately minimal and does not resolve all failure cases, it demonstrates the potential of incorporating visual feedback loops for execution robustness. Future extensions may incorporate more sophisticated recovery strategies, such as policy rollback, re-grasping, or online subgoal regeneration, to further enhance success rates under real-world uncertainties.
3.3 Long-horizon Manipulation Analysis
We analyze performance on LIBERO-Long, which features multi-step tasks with sequential subgoals. Table 2 reports subgoal-level and overall success rates across 10 manipulation scenarios. Detailed task descriptions are provided in Appendix D.
| Model | Soup-Sauce | Cheese-Butter | Stove-Moka | Bowl-Drawer | Mug-Mug | ||||||||||
| Soup | Sauce | SR | Cheese | Butter | SR | Stove | Moka | SR | Bowl | Drawer | SR | Mug | Mug | SR | |
| OpenVLA | 0.72 | 0.52 | 0.46 | 0.86 | 0.64 | 0.64 | 0.88 | 0.64 | 0.64 | 0.44 | 0.32 | 0.32 | 0.60 | 0.58 | 0.44 |
| Agentic Robot | 0.83 | 0.67 | 0.67 | 0.88 | 0.78 | 0.78 | 0.88 | 0.71 | 0.71 | 0.72 | 0.56 | 0.56 | 0.71 | 0.63 | 0.63 |
| Model | Book-Caddy | Mug-Pudding | Soup-Cheese | Moka-Moka | Mug-Wave | ||||||||||
| Book | Caddy | SR | Mug | Pudding | SR | Soup | Cheese | SR | Moka | Moka | SR | Mug | Wave | SR | |
| OpenVLA | 0.98 | 0.82 | 0.82 | 0.64 | 0.54 | 0.54 | 0.72 | 0.64 | 0.60 | 0.58 | 0.22 | 0.22 | 0.54 | 0.50 | 0.46 |
| Agentic Robot | 0.98 | 0.84 | 0.84 | 0.65 | 0.60 | 0.60 | 0.83 | 0.64 | 0.64 | 0.64 | 0.17 | 0.17 | 0.66 | 0.72 | 0.58 |
Performance gains. The Agentic Robot consistently outperforms OpenVLA, with an average improvement of 12.1% across all tasks. The most significant gains are observed in challenging scenarios that have lower baseline success rates, such as the following: in the Bowl-Drawer task, there is a 24% improvement; in the Mug-Mug task, a 19% improvement; and in the Soup-Sauce task, a 21% improvement. These results confirm the effectiveness of our approach in reducing cascading failures.
Error prevention. OpenVLA frequently propagates errors through subtasks by continuing even when execution is incomplete. In contrast, Agentic Robot employs VLM-based verification, which ensures that progression cannot occur until each subgoal is confirmed as complete. This approach accounts for the improvements observed in spatially complex tasks like Stove-Moka, which saw a 7% increase in performance, highlighting the importance of robust checkpoint verification.
Limitations. In the Moka-Moka task, the Agentic Robot demonstrates decreased performance in one subgoal (17% vs. 22%), highlighting a limitation in managing scenes that require fine-grained coordination for the placement of symmetric objects. A qualitative analysis shows that while the first Moka Pot is placed correctly, the system often centers it on the stove surface. This positioning leaves insufficient space for the second Moka Pot, resulting in a placement failure despite having a correct high-level plan. This situation underscores the current weakness of the Agentic Robot in anticipating spatial constraints and resolving conflicts between similar subgoals that involve identical object types. Future work may address this issue by implementing temporal structure modeling, spatial intent prediction, or memory-aware policies that explicitly consider previous placements during planning and execution.
3.4 Verification Frequency Analysis
We investigate how verification frequency affects both performance and computational efficiency by evaluating Agentic Robot with verification every 10, 20, or 50 steps. Fig. 4 presents the results.
Success rate sensitivity. All verification frequencies yield similar success rates for the Spatial, Object, and Goal suites, indicating that these shorter tasks are resilient to verification sparsity due to their limited subgoal durations and lower risk of error propagation. In contrast, LIBERO-Long shows considerable sensitivity to verification frequency. When verification is reduced to every 50 steps, the success rate decreases by 6 percentage points (from 61.8% to 55.8%). Providing verification every 10 steps does not offer any additional benefits beyond the 20-step interval. This suggests that long-horizon tasks require frequent validation to prevent error accumulation, but the benefits diminish beyond an optimal verification frequency.
Efficiency-Performance trade-off. Verification frequency substantially impacts execution time. High-frequency verification (every 10 steps) increases episode duration due to VLM inference overhead, particularly in LIBERO-Long, where the runtime difference between 10-step and 50-step intervals exceeds 15 seconds per episode.
Optimal configuration. Our analysis identifies the 20-step interval as optimal: it maintains peak success rates while reducing computational overhead. We adopt this as our default configuration. These findings suggest that while frequent verification is critical for long-horizon robustness, it introduces unnecessary computational cost for simpler tasks. Future work could explore adaptive verification schemes that dynamically adjust frequency based on task complexity, execution uncertainty, or environmental dynamics.
3.5 Ablation Study
To analyze each component’s contribution, we conduct a systematic ablation study on LIBERO-Long, where extended manipulation sequences are particularly sensitive to architectural modifications. Table 3 summarizes our findings.
Multimodal Planning. When the planner is restricted to text-only inputs, the task success rate drops to 57.4%, representing a 4.4% decline relative to the full multimodal setting. This performance gap highlights the importance of visual context for grounding instructions and resolving object ambiguity, particularly in cluttered scenes.
Recovery Mechanism. Removing the recovery routine after failed subgoal verification leads to a reduced success rate of 59.7%, reflecting a 2.1% degradation. This result confirms the value of even minimal corrective behaviors in mitigating error accumulation across long-horizon task sequences.
| Setting | SR (%) |
| No Visual Input | 57.4 |
| No Recovery Mechanism | 59.7 |
| No Fine-tuned VLM | 35.3 |
| No Subgoal Decomposition | 53.7 |
| Full System | 61.8 |
Verification Quality. Substituting the fine-tuned verifier with a zero-shot VLM yields a substantial drop in success rate to 35.3%, indicating a degradation of 26.5%. This sharp decline suggests that generic models are insufficiently sensitive to subtle changes in scene state, and that domain adaptation is essential for accurate subgoal-level introspection. As illustrated in Fig. 5, the zero-shot verifier fails to detect subgoal completion, causing repetitive execution and eventual failure. In contrast, our fine-tuned model enables sequential verification and completion of each subtask. Quantitatively, it improves the LIBERO-Long success rate by +7.8% over the zero-shot baseline, confirming the value of even small-scale domain adaptation.
Hierarchical structure. Without subgoal decomposition (i.e., vanilla OpenVLA), requiring the system to perform instructions as single, monolithic goals results in a performance drop to 53.7%, a decrease of 8.1%. This supports our main hypothesis that breaking complex tasks into smaller, atomic skills significantly enhances execution reliability and verification accuracy. As illustrated in Fig. 6, the lack of subgoal decomposition causes the executor to attempt the full task without the ability to recover from intermediate failures, ultimately skipping necessary steps and leading to task failure. In contrast, our hierarchical agent sequentially verifies and completes each subtask, ensuring successful full-task completion. Each component offers measurable benefits, with fine-tuned verification and hierarchical planning contributing the most substantial improvements. These findings confirm the effectiveness of our Agentic Robot for handling long-horizon manipulation tasks.
4 Discussion and Limitations
We now reflect on the design of Agentic Robot, highlighting its strengths and pointing out remaining challenges.
Verification as a robustness mechanism. A central contribution of our framework is the introduction of visual verification as a control signal for subgoal progression. The verifier functions as a semantic gatekeeper that determines whether to proceed, retry, or terminate, enabling subgoal-level error detection and correction without access to ground-truth state information. Our empirical results demonstrate this approach’s effectiveness in mitigating compounding errors, particularly in long-horizon settings where early mistakes could cascade through subsequent action sequences. The incorporation of recovery behaviors further enhances system resilience under environmental uncertainty and partial observability.
Real-World deployment challenges. While our results are validated in high-fidelity simulated environments, transferring Agentic Robot to physical platforms introduces several challenges. These include handling sensor noise in RGB inputs, adapting to real-world lighting variations and occlusions, and compensating for actuation delays. Furthermore, the verifier’s robustness to visual domain shift requires extensive evaluation. Future work will incorporate domain adaptation and sim-to-real transfer techniques, particularly focusing on real-image fine-tuning for both the verifier and executor components to mitigate these challenges.
Adaptive verification scheduling. Currently, verification occurs at fixed intervals (every 20 frames), regardless of task complexity, execution speed, or object dynamics. Although effective in our evaluations, this heuristic approach is likely suboptimal for computational efficiency. We propose to explore adaptive verification strategies that leverage confidence-aware scheduling based on motion intensity, subgoal typology, or the LLM’s uncertainty quantification. Such approaches would optimize computational resource allocation while maintaining task safety and correctness guarantees.
5 Conclusion
This work introduces Agentic Robot, a brain-inspired framework that uses Standardized Action Procedure (SAP) to improve reliability and interpretability in robotic manipulation systems. The framework decomposes complex tasks into coordinated interactions between three specialized components–planner, executor, and verifier–operating through well-defined SAP protocols that mirror biological cognition. By establishing explicit protocols for component communication, progress monitoring, and failure recovery, SAP addresses fundamental limitations in existing manipulation systems while allowing independent component optimization through standardized interfaces. Extensive validation on the LIBERO benchmark demonstrates state-of-the-art performance with 79.6% average success rate, including substantial improvements of 24% on Bowl-Drawer tasks and 21% on Soup-Sauce tasks. This successful integration of brain-inspired architectures with standardized coordination protocols shows how biologically motivated design principles can enhance both performance and interpretability in embodied AI systems.
Appendix of paper “Agentic Robot: A Brain-Inspired Framework for Vision-Language-Action Models in Embodied Agents”
Appendix A Detailed Experiment Setup
A.1 Benchmark
Our primary evaluation is conducted on the LIBERO suite, a benchmark collection for instruction-following manipulation in diverse simulated environments. We select the four suitesLIBERO-Long, LIBERO-Spatial, LIBERO-Object, and LIBERO-Goaleach comprising 10 tasks and 50 human-teleoperated demonstrations per task. The multitask performance of the pretrained VLA policy is evaluated on these suites. Specifically:
LIBERO-Spatial: Contains the same set of objects but in varying layouts, testing the model’s ability to understand spatial relationships. Example language instruction: pick up the black bowl in the top drawer of the wooden cabinet and place it on the plate.
LIBERO-Object: Features consistent scene layouts but introduces different objects, evaluating the model’s understanding of object types. Example language instruction: pick up the chocolate pudding and place it in the basket.
LIBERO-Goal: Maintains the same objects and layouts while varying task goals, assessing the model’s knowledge of diverse task-oriented behaviors. Example language instruction: turn on the stove.
LIBERO-Long (also referred to as LIBERO-10): Comprises long-horizon tasks involving diverse objects, layouts, and task goals, challenging the model’s ability to handle extended planning and execution. Example language instruction: pick up the book and place it in the back compartment of the caddy.
A.2 Baselines
Baselines. We benchmark our approach against the following generalist policies, including state-of-the-art open-sourced models:
Diffusion Policy [5]: A way of generating robot behavior by representing a robot’s visuomotor policy as a conditional denoising diffusion process.
TraceVLA [38]: Finetuned from OpenVLA with visual trace prompting.
SpatialVLA [23]: A 4B parameter VLA trained on 1.1 million real-world robot episodes.
Appendix B Related Works
B.1 Large Reasoning Models
Recent progress in large reasoning models (LRMs) has dramatically improved general-purpose cognitive capabilities, providing a foundation for downstream embodied agents [29]. Models such as DeepSeek-R1 [8] adapt rule-based reinforcement learning pipelines into massive 671B-parameter text models, enabling distilled checkpoints for multimodal adaptation. Gemini-2.5 [27] unifies vision, audio, and long-text inputs into a state-of-the-art retrieval system over 10M-token contexts. MM-Eureka [20] introduces visual-math chain-of-thought training with reinforcement learning, achieving new benchmarks on multimodal math reasoning. Similarly, VLM-R1 [24] transfers the R1 architecture into the vision-language domain, yielding strong zero-shot visual reasoning capabilities. Although these models offer unprecedented perceptual and reasoning skills, they are not explicitly trained for robotics tasks, leaving open the challenge of grounding abstract reasoning into actionable physical plans, particularly in long-horizon settings.
B.2 Vision-Language Models in Robotics
Vision-language models (VLMs) have increasingly been adapted into robotic systems to bridge perception and semantic understanding [37]. PaLM-E [6] integrates vision tokens into a 562B parameter LLM, enabling long-horizon real-world manipulation while retaining general VQA abilities. VoxPoser [12] leverages CLIP and LLM prompting to synthesize volumetric value maps for zero-shot pick-and-place. In navigation, VLM-Social-Nav [25] grades candidate trajectories through captioning models to enhance social compliance, while NaVid [36] predicts step-wise actions from videos and language for map-free instruction following. GSON [19] further extends visual reasoning to group-aware navigation by detecting social formations. While these systems demonstrate the semantic reasoning power of VLMs, they primarily focus on static goal conditions or trajectory scoring, rather than dynamically verifying subgoal progress within a continuous manipulation task as required by long-horizon execution.
B.3 Vision-Language-Action Models
Vision-Language-Action (VLA) models directly map visual observations and language instructions to robotic control actions [34, 18, 30]. Notable examples include RT-2 [41], which co-trains a VLM with robot episodes using action tokenization, and OpenVLA [14], which scales this approach with 970k demonstrations, outperforming RT-2-XL while maintaining efficiency. Advances like ChatVLA [39] incorporate phased alignment and policy routing to preserve reasoning capabilities during execution, while ECoT [35] introduces chain-of-thought reasoning for improved performance on long-horizon tasks. RoboMM [33] achieves cross-domain generalization through modality-isolation masking across diverse datasets. SpatialVLA [23] enhances 3D spatial understanding with Ego3D position encoding and adaptive action grids, while TraceVLA [38] improves spatial-temporal awareness by visualizing robot trajectory traces. Despite these advances, current VLA systems typically lack mechanisms for subgoal monitoring and dynamic recovery, limiting their robustness in complex, multi-stage environments.
Appendix C Atom Skill Library
An atomic skill library is a dynamically expanding repository of fine-grained manipulation skills that a robot can invoke to perform complex end-to-end tasks without retraining a monolithic policy [15]. In our framework, the atomic skill library is a curated collection of low-level manipulation primitives, such as “pick up [object],” that have been manually defined by domain experts to guarantee predictable performance and semantic clarity. Each atomic skill encapsulates a self-contained control policy or trajectory generator, allowing complex tasks to be decomposed into a sequence of reusable, verifiable subroutines. Although hand-crafting these primitives ensures reliability and interpretability, the same library structure can be populated or even expanded automatically by large language models: an LLM can translate high-level task descriptions into candidate skill definitions or suggest refinements to existing primitives. By combining expert-driven design with LLM-powered generation, this hybrid approach accelerates skill coverage, simplifies adaptation to novel tasks, and preserves the modularity and transparency crucial for robust robotic operation. The atom skill library is detailed as follows:
pick up [object] (from [location]/[object])
place [object] in/on [location]/[object]
push [object] to [location]
place [object] to the [direction] of [object]
open/close [object/container/drawer/etc.]
turn on/off [device]
Appendix D Experiments on Task Division
D.1 Description of LIBERO-Long
Now, we give the detailed task description of LIBERO-Long as used in Section 3.3.
Soup-Sause: The robot must locate and pick up both the alphabet soup can and the tomato sauce can. It then needs to place both items inside a basket. This task tests multi-object handling and proper placement.
Cheese-Butter: In this task, the agent must pick up two items: a box of cream cheese and a piece of butter. Both need to be placed into the same basket. The challenge involves identifying similar-looking food items and executing sequential pick-and-place actions.
Stove-Moka: The robot is required to first turn on a stove, then place a moka pot on top of it. This involves both environment interaction (activating the stove) and precise object placement. It tests sequential decision-making and tool use.
Bowl-Drawer: The task requires the agent to open a bottom drawer of a kitchen cabinet, place a black bowl inside, and then close the drawer. This combines manipulation of articulated components (the drawer) with careful object handling.
Mug-Mug: The agent needs to distinguish two mugs by color and place them on specific plates: the white mug on the left plate, and the yellow-and-white mug on the right. The task emphasizes color-based object recognition and spatial arrangement.
Book-Caddy: The robot must pick up a book and place it into the rear compartment of a caddy organizer. This task involves handling flat objects and placing them into confined spaces, testing precision and spatial reasoning.
Mug-Pudding: The robot places a white mug onto a plate and then positions a chocolate pudding to the right of that plate. It requires understanding relative spatial positioning between objects and accurate placement.
Soup-Cheese: This task is similar to task 1 but with a different object combination: alphabet soup and cream cheese box. The agent must place both items into a basket, reinforcing generalization across similar multi-object tasks.
Moka-Moka: The robot needs to find two moka pots and place both on the stove. The challenge lies in handling duplicate objects and placing them correctly on the same surface.
Mug-Wave: In this task, the agent must place a yellow-and-white mug inside a microwave and then close the microwave door. It tests interaction with articulated appliances and precise object insertion.
D.2 Illustration of Task Division
Appendix E Pseudo Code of Agentic Robot
Appendix F Pseudo Code for Perception-based Verifier
Appendix G Prompt of LRM for Zero Shot Task Devision